OpenStudy (anonymous):

What is -2≤d+1/2<1/2 = ?

OpenStudy (anonymous):

it is not equal to anything i bet the problem is solve \[-2\leq d +\frac{1}{2}<\frac{1}{2}\] for \(d\)

OpenStudy (anonymous):

solve in one step by subtracting \(\frac{1}{2}\) straight across

OpenStudy (anonymous):

-2≤d<0

OpenStudy (anonymous):

no, left hand side should be \(-2.5\) or \(-\frac{5}{2}\)

OpenStudy (anonymous):

thanks for your replies guys :)